A crystal of MgCl2 contains 5.85 x 10e20 cations.
a) how many formula units are in the crystal?
b) how many moles of MgCl2 are in the crystal?
c) what is the mass of the crystal?
1
Expert's answer
2019-10-25T05:30:53-0400

a) As 1 formula unit of MgCl2 contains ine ion of Mg2+ then

"N(Mg^{2+}) = N(MgCl_2)= 5.85\times10^{20} ions"

b)

"n(MgCl_2)=\frac{N}{N_A} = \frac{5.85 \times10^{20}}{6.022\times10^{23}} = 9.71\times 10^{-24} mol"

c)

"m(MgCl_2)= n\times M = 9.71\times10^{-4}\times95.21 = 9.25\times10^{-2}g"